OpenStudy (anonymous):

The perimeter of a rectangular garden is 40 meters. The length of the garden is 1 meter less than twice its width. What is the length of the garden? 7 meters 10 meters 13 meters 16 meters

OpenStudy (angela210793):

length=x and width=y we have 2x+2y=40m also x=2y-1

OpenStudy (anonymous):

hi angela...

OpenStudy (angela210793):

hey :)

OpenStudy (anonymous):

how are u?

OpenStudy (angela210793):

so 40=2(2y-1)+2y=6y-2 y=7 so x=2*7-1=13

OpenStudy (angela210793):

Fine thnx,How are you?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

good...so u r sleep is ok..and u r fit for ur tests? good to hear tat..i m just fine now..thnx for asking:)

OpenStudy (angela210793):

yea I'm much better....(actually have started to drink coffee :P) and i learnt 5 authors 2day :) Shakespeare,camus,kafka,balzac,and flober

OpenStudy (anonymous):

wow! ur great! keep up..but dont take too much coffee..it harms ur nervous sytem ..:)

OpenStudy (angela210793):

nah...just a very small cup and lots of sugar :):):) Mom's calling 4 dinner..I'll b back in a few mins :):)

OpenStudy (anonymous):

:) take care..

OpenStudy (angela210793):

I'm back :):) What have u been doing?

OpenStudy (anonymous):

well..have been helping someone..:) do u have any problems for me..u had 300..remember? :)

OpenStudy (angela210793):

yea i remember..how can 300 problems b forgotten? :P i'm dealing with literature now so...no time for them...I'll ask ur help after june 1st ;) :)

OpenStudy (anonymous):

k..waiting anxiously till then :P ..

OpenStudy (angela210793):

lol :P:P:P I'll give them to you(if i'm not dead till june 1st) :P

OpenStudy (anonymous):

hope u succeed:) best of luck with ur exams..(esp. literature..)

OpenStudy (angela210793):

^_^ THNX :):):)